TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Created by Main.MirkoPojer on 2007-05-25 - 21:28

Powering Tests in Sector 7-8: activity of Friday 25/05.

The day was fully devoted to perform tests on D2.

The following tests have been executed on D2 (MANUAL execution, due to change in procedure):
  • PIC2 (SEQUENCER execution); test successful;
  • PLI1: ramp to 350 A with 10 A/s ramp rate, we stayed on the plateau 4 min and then provoked a FPA; test successful;
  • PLI2: ramp to 350 A with 10 A/s ramp rate, then ramp to 1000 A with 10 A/s, and quench driven by QPS, generating an artificial quench signal from one current lead detector; the test was successful;
  • PNO: during the first attempt to ramp to 3000 A, the power converter went to ‘faults’ and produced a FPA when reaching 1008 A; investigations of PO showed that a hardware limitation was present. The second attempt was successful:
    • Ramp to 350 A with 10 A/s ramp rate;
    • Continue ramping to 3000 A; plateau of 4 min;
    • Continue ramping to 6000 A (nominal current); plateau of 30 min to check voltage drift;
    • Provoked quench at nominal current.

Proposal for the activity of next week.

  • Continue powering Q4 and D2
  • Continue powering the correctors (4 circuits have already passed the first 2 steps, three more are available).
  • Possible powering of the whole matching section.
  • Powering of the arc 60 A orbit correctors, according to humidity conditions.

  • Continue installation of plastic bags on DFBAN and DFBMH.
  • Check the status of condensation of the 13kA leads of DFBAO which have copper laminations and no plastic bags; check the status of the 6kA lead of DFBMA which has copper laminations and no plastic bags.
  • Installation of ventilators on the 60 A orbit correctors leads.

  • ElQA on the selected circuits: see presentation of Davide in the last post.

  • Quench heaters qualification, without interference with ElQA.

Edit, attach file or add comment

Created by Main.MirkoPojer on 2007-05-25 - 21:15

Powering Tests of Sector 7-8: minutes of the meetings of 25/05.

PDF Q4.L8.pdf:

PDF RAT_20070525.pdf:

PDF ELQA_CCC_25052007.pdf:


Edit, attach file or add comment

Created by Main.BorisBellesia on 2007-05-25 - 11:33

Short Circuit Tests in UA27 - 25th May

PDF RAT_20070525.pdf:

EE - IST    
13kA EE DONE 16-04-07
600A EE DONE 23-04-07
Circuits >2kA    
Powering tests PC 4-8kA DONE 11-05-07
8h run 4-8kA - 29-05-07
Interlock tests RQD&F ONGOING 02-05-07
Powering tests RQD&F - 24-05-07
Preparation RB.A23 ON GOING 19-04-07
Interlock tests RB.A23 DONE 25-05-07
Powering tests RB.A23 - 30-05-07
8h run >= 8kA - 04-06-07
Circuits < 2kA    
120A powering tests DONE 03-05-07
600ANOEE powering tests DONE 08-05-07
600A-EE interlock tests DONE 03-05-07
600A-EE powering tests DONE 18-05-07
8 hr run PC < 2kA - 29-05-07
24hr heat run test - 07-06-07
Polarity Check < 8KA - 30-05-07
AUG- UPS test - 08-06-07

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-05-02 - JacekSzkutnik
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login